# A Gross-Pitaevskii theory for an excitonic incompressible Bose solid

### v1

@@ Abstract
For many years there has been considerable theoretical interest in the Bose solid, but it has never been definitively observed experimentally. A true Bose solid must have a lattice site occupancy of exactly one boson per site [1]. Precisely one boson per site generates a particle-antiparticle gap in the energy spectrum, and as a consequence, such a Bose solid will be incompressible. References [2, 3] showed that the ground state of a Bose solid is an incompressible supersolid with the particles localized on their own lattice sites by strong particle-particle interactions, while at the same time they are in a single quantum condensate. The Bose solid/incompressible supersolid is quite different from the superfluids with periodic clustering [4, 5] or density modulation [6] in dipolar cold atom systems that have been termed supersolid-like.
In this letter, we generalize the Gross-Pitaevskii approach to map out the ground state of the Bose solid at zero temperature. The Gross-Pitaevskii formalism is mean-field Hartree, so it includes interactions of a particle with itself [7, 8]. However in a Bose solid with only a single particle on each site, there should be no self-interaction terms [9]. Here, we extend the conventional Gross-Pitaevskii approach to capture this property. Furthermore, exchanges of bosons in the Bose solid may take place with difficulty because of the presence of other neighbors [1]. We therefore also incorporate in our Gross-Pitaevskii approach the effect of strong short-range two-particle correlations.
Our results show that excitons in a double-layer semiconductor heterostructure can form a Bose solid and would constitute the first experimental realization of an incompressible supersolid. In this system, the excitons are spatially indirect, formed of electrons and holes that are confined in two adjacent two-dimensional conducting layers separated by a thin insulating barrier [10, 11]. With equal electron and hole densities, the exciton dipoles will be aligned perpendicular to the layers so the dipole-like exciton interactions are purely repulsive. Sufficiently strong repulsion can localize the excitons on separate lattice sites of a dipolar crystal [12, 13] similarly to electrons in a Wigner crystal [14]. This state differs from cold atom systems where attractive as well as repulsive atomic interactions acting in competition are needed to generate periodic clusters [15] which in general are not incompressible. We have proposed the existence of an exciton incompressible supersolid, based on a variational calculation [16].
The Hamiltonian for the excitonic state is
@@
We take equal electron and hole densities, the exciton dipoles align perpendicular to the layers, and so the exciton-exciton interactions are purely repulsive,
@@
^$\widehat{{\cal{H}}}^{\rm{eff}}=\int$ $d^{2}\mathbf{r}$ $\,\widehat{\psi}^{\dagger}(\mathbf{r})\left(-\frac{\hbar^{2}\nabla^{2}}{2M_{\text{X}}}\right)\widehat{\psi}(\mathbf{r})$
@@
For sufficiently strong exciton interactions, a normal-state exciton crystal can form, with exactly one exciton per lattice site [12, 13]. This normal-state Bose solid can be described by a many-body wave function consisting of a product of single-particle wave functions $\phi_{i}(\mathbf{r})$, each $\phi_{i}(\mathbf{r})$ centered on a different site $i$ of the periodic lattice [2]. We derive a Hartree equation for this state in the supplemental material (SM) [18], as
@@
The attractive interaction term in Eq. (21) is generated from the property that, while an exciton on a singly-occupied site $i$ feels the repulsive interactions from the excitons in the neighboring unit cells, it does not feel its own potential within the unit cell $i$ [1] (see Eq. (8) in the SM [18]). This term removes the contributions in the Hartree equation from a single boson in a unit cell interacting with itself. The term is only present when there is exactly one particle per unit cell, since with more particles there would be interactions within the unit cell. The effect of the absence of these self-interactions within a site acts like a potential well centered on the site. For sufficiently strong repulsion between the excitons, the potential wells can become deep enough for bound-state solutions to exist [1]. In fact, in systems with purely repulsive particle-particle interactions such as electrons or perpendicularly aligned excitons, this is the mechanism that can drive the solidification [2], and in these repulsive systems quantum Monte Carlo simulations have confirmed the existence of normal solids [14, 12, 13]. The solid would be incompressible since the potential wells lead to a prohibitive energy cost against doubly occupied sites [9], analogous to a Mott insulator.
@@
In the present study we investigate the possibility of solidification coexisting with quantum coherence in a supersolid. We extend the standard Gross-Pitaevskii formalism to include the effect of the absence of self interactions, specific to this exciton system. In place of the many-body wave function of the normal state system, we construct the order parameter for the incompressible supersolid of the form, $\Psi(\mathbf{r})=\sum_{i=1}^{N}\phi_{i}(\mathbf{r})$, where $i$ is the site index for the $N$ sites. With this construction, each exciton has the same wave function spread across the entire lattice and with peaks at every site. This order parameter results in both off-diagonal long-range order and also diagonal long-range order coming from the lattice symmetry. In addition, for a supersolid there must exist phase locking between adjacent sites. This arises from mutual overlap of the $\phi_{i}(\mathbf{r})$ [19]. If the $\phi_{i}(\mathbf{r})$ were exponentially localized, it becomes impossible to link the phases on the different lattice sites, since thermal and zero-point quantum fluctuations would eliminate any coherence between the near-negligible overlaps. In order to monitor the overlaps, we introduce a parametrization of the order parameter. We represent $\phi_{i}$ and $\Psi$ in the form,
$\phi_{i}(\mathbf{r})$ $=$ $\frac{\sqrt{\rho_{B}}}{N}+G(\mathbf{r}-\mathbf{r}_{i})\ ,$ (8)
$\Psi(\mathbf{r})$ $=$ $\frac{1}{\sqrt{N}}\left(\sqrt{\rho_{B}}+\sum_{i=1}^{N}G(\mathbf{r}-\mathbf{r}_{i})\right)\ .$ (9)
This measures the overlap by means of a uniform background component $\rho_{B}$. In cases when $\rho_{B}$ tends to zero there is negligible overlap so there would be no supersolid. The $G(\mathbf{r}-\mathbf{r}_{i})$ are functions localized on the different sites $i$ and centered at $\mathbf{r}_{i}$. In cases when the $G(\mathbf{r}-\mathbf{r}_{i})$ reduce to zero, Eq. (24) represents the homogeneous order parameter for a superfluid, $|\Psi(\mathbf{r})|^{2}\sim\rho_{B}$.
@@
$\frac{\hbar^{2}\nabla^{2}}{2M_{\text{X}}}\Psi(\mathbf{r})+U_{\text{XX}}(\mathbf{r})\Psi(\mathbf{r})=\mu\Psi(\mathbf{r})\ ,$ (10)
with mean-field potential,
$U_{\text{XX}}(\mathbf{r})$ $=$ $\int d^{2}\mathbf{r}^{\prime}\,t_{\text{XX}}(|\mathbf{r}^{\prime}-\mathbf{r}|)|\Psi(\mathbf{r}^{\prime})|^{2}$ (11)
$-\int_{\mathbf{r}^{\prime}\in i_{\mathbf{r}}}\!\!\!\!\!\!d^{2}\mathbf{r}^{\prime}t_{\text{XX}}(|\mathbf{r}^{\prime}-\mathbf{r}|)\,|\Psi(\mathbf{r}^{\prime})|^{2}\ .$
@@
The potential $U_{XX}(\mathbf{r})$ differs from the potential in the standard Gross-Pitaevskii equation for dipolar bosons [21], not only because $V_{\text{XX}}$ has been replaced by $t_{\text{XX}}$ to account for two-body correlations, but also because of the additional potential term. It is this term, that accounts for the absence of self-interactions, which makes supersolid solutions possible.
@@
We solve Eqs. (10) and (11) by evolving in imaginary time [22], starting from a spatially randomized order parameter. Note throughout the numerical calculation there is no restriction that $\Psi(\mathbf{r})$ is of the form of Eq. (24). Because of the non-local interaction $V_{\text{XX}}(\mathbf{r})$, we evaluate Eq. (11) by means of the convolution theorem [23]. For simplicity, we define the $i_{\mathbf{r}}$ in Eq. (11) on a square lattice. For exciton density $\rho$, the lattice constant is given by $a=2/(\pi\sqrt{\rho})$ for one exciton per site. A triangular lattice may have a slightly lower energy, but we leave the issue of square vs. triangular lattice outside the scope of the present study. All simulations were performed on a $256^{2}$ grid covering the sample spatial area. Solutions were taken as converged when the relative change in energy in consecutive time steps decreases below $10^{-8}$. After reaching convergence, we have verified for self-consistency that $\Psi(\mathbf{r})$ indeed has the form of Eq. (24), with $\rho_{B}$ not zero.
We focus first on the low-density regime, setting the average interparticle spacing $r_{0}=1/(\pi\sqrt{\rho})=30$. As length and energy scales, we use the effective Bohr radius $a_{B}=\hbar^{2}\epsilon/e^{2}m_{e}^{*}$ and the effective Rydberg Ry. Figure 1(a) shows the ground-state order parameter $|\Psi(\mathbf{r})|^{2}$ for small interlayer separation $d=1$. The solution $\Psi(\mathbf{r})$ is homogeneous – corresponding to a superfluid: for such a small $d$, the exciton dipole interactions are too weak to drive the system to a solid.
@@
Figures 1(c) and 1(d) show that when the interlayer spacing $d$ is further increased, the peak in $\Psi(\mathbf{r})$ on each site narrows and the amplitude $\rho_{B}$ of the homogeneous background component decreases.
@@
Figure 2(b) compares the supersolid energy eigenvalue $\mu_{SS}$ from Eq. (10) determined using $U_{XX}(\mathbf{r})$ from Eq. (11) and the superfluid energy eigenvalue $\mu_{SF}$ determined without the subtracted self-interaction term in $U_{XX}(\mathbf{r})$. For $d<6$ the two energy eigenvalues are identical, indicating a stable superfluid, but for $d\gtrsim 6$ the energy eigenvalue $\mu_{SF}$ is higher than $\mu_{SS}$ indicating that the incompressible supersolid becomes the ground state for $d\gtrsim 6$.
@@
Figure 2: (a) Dependence on interlayer separation $d$ of $\rho_{B}/\rho$, the amplitude of the homogeneous background component of $\Psi(\mathbf{r})$, for $r_{0}=30$. (b) Energy eigenvalue of the supersolid state (SS) and corresponding eigenvalue of the superfluid state (SF), calculated without the self-interaction correction.
@@
Figure 3: The phase diagram at zero temperature, as a function of interlayer separation $d$ and average interparticle spacing $r_{0}$. Dashed red line: superfluid to supersolid transition. Dash-dotted grey line: condensate collapse to a normal liquid [29]. Dashed green line: exciton liquid to normal solid transition from Ref. [12]. Dotted black line: $d=r_{0}$. Insets: order parameter profile $|\Psi(\mathbf{r})|^{2}$ at the points indicated by grey dots.
Within the supersolid phase, as the density is increased the exciton repulsive interaction (Eq. (2)) evolves from dipolar-like to Coulomb-like, making the supersolid soften and eventually melt. Just before melting, the system enters the BEC-BCS crossover regime where screening rapidly becomes strong and the system simultaneously loses coherence and crystal structure. The quantum condensed phases lie almost entirely within the BEC regime where screening is negligible, allowing us to describe the system in terms of a Bose condensate with a macroscopic order parameter. This makes our Gross-Pitaevskii formalism valid across the area of the phase diagram encompassing the supersolid and superfluid phases.
@@
Acknowledgements
We thank Filippo Pascucci for very useful discussions. This work was supported by Research Foundation – Flanders (FWO-Vl), with contract number 12A3T24N; by the University Research Fund (BOF), with project numbers 53228, G0A9F25N, G0AIY25N, GOH1122N, G061820N, and G060820N; and by the Brazilian Council for Research (CNPq), through the PRONEX/FUNCAP, Universal, and PQ programs.

### v2

@@ Abstract
For many years there has been considerable theoretical interest in the Bose solid, but it has never been definitively observed experimentally. A true Bose solid must have a lattice site occupancy of exactly one boson per site [1]. Precisely one boson per site generates a particle-antiparticle gap in the energy spectrum, and as a consequence, such a Bose solid will be incompressible. References [2, 3] showed that the ground state of a Bose solid is an incompressible supersolid with the particles localized on their own lattice sites by strong particle-particle interactions, while at the same time they are in a single quantum condensate. The Bose solid/incompressible supersolid is quite different from the superfluids with periodic clustering [4, 5] or density modulation [6] in dipolar cold atom systems that have been termed supersolid-like. We have proposed the existence of such an incompressible supersolid in an exciton system based on a variational calculation [7].
In this letter, we generalize the Gross-Pitaevskii approach to map out the ground state of the Bose solid at zero temperature. The Gross-Pitaevskii formalism is mean-field Hartree, so it includes interactions of a particle with itself [8, 9]. However in a Bose solid with only a single particle on each site, there should be no self-interaction terms [10]. Here, we extend the standard Gross-Pitaevskii approach to capture this property. Furthermore, exchanges of bosons in the Bose solid may take place with difficulty because of the presence of other neighbors [1]. We therefore also incorporate in our Gross-Pitaevskii formalism the effect of strong short-range two-particle correlations. Development of a Gross-Pitaevskii based approach is a crucial enabler for further studies of static and dynamic phenomena in exciton superfluids and supersolids.
Our results show that excitons in a double-layer semiconductor heterostructure can form a Bose solid and would constitute the first experimental realization of an incompressible supersolid. In this system, the excitons are spatially indirect, formed of electrons and holes that are confined in two adjacent two-dimensional conducting layers separated by a thin insulating barrier [11, 12]. With equal electron and hole densities, the electric field from the gates aligns the exciton dipoles perpendicular to the layers so they cannot deform or rotate. This ensures the dipole-like exciton interactions always purely repulsive, in contrast to free dipoles. Sufficiently strong repulsion can localize single excitons on separate lattice sites of a dipolar crystal [13, 14] similarly to electrons in a Wigner crystal [15]. In both cases the purely repulsive interaction ensures only one particle per site. This differs from cold atom systems where attractive as well as repulsive atomic interactions acting in competition are needed to generate periodic clusters [16] which in general are not incompressible.
We start with the Hamiltonian for the excitonic state,
@@
The area of the confined system is fixed by gates. The gates also fix the density and the number of particles. We take equal electron and hole densities, the exciton dipoles align perpendicular to the layers, and, so the repulsive exciton-exciton interactions are,
@@
$\widehat{{\cal{H}}}^{\rm{eff}}=\int$ $d^{2}\mathbf{r}$ $\,\widehat{\psi}^{\dagger}(\mathbf{r})\left(-\frac{\hbar^{2}\nabla^{2}}{2M_{\text{X}}}\right)\widehat{\psi}(\mathbf{r})$
@@
For sufficiently strong exciton interactions, a normal-state exciton crystal can form, with exactly one exciton per lattice site [13, 14]. This normal-state Bose solid can be described by a many-body wave function consisting of a product of single-particle wave functions. However in the absence of symmetry breaking, the solid many-body wave function will be translationally invariant, consisting of a superposition of an infinite number of infinitesimally translationally-shifted degenerate many-body states. The translation symmetry is broken by selecting one of these degenerate states, thereby making a choice of a particular set of sites $\{a_{i}\}$ in a fixed periodic lattice. The single-particle wave functions $\phi_{i}(\mathbf{r})$ will then be localized on the different sites $i$ of the lattice [2]. We derive a Hartree equation for this state in the supplemental material (SM) [18], as
@@
The effective attractive interaction term in Eq. (19) is generated from the property that, while an exciton on a singly-occupied site $i$ feels the repulsive interactions from the excitons in the neighboring unit cells, it does not feel its own potential within the unit cell $i$ [1] (see Eq. (8) in the SM [18]). This term removes the contributions in the Hartree equation from a single boson in a unit cell interacting with itself. The term is only present when there is exactly one particle per unit cell, since with more particles there would be interactions within the unit cell. This term is not an externally imposed potential but self-consistently arises from the exciton-exciton interactions.
The effect of the absence of these self-interactions within a site acts like a potential well centered on the site. For sufficiently strong repulsion between the excitons, the potential wells can become deep enough for bound-state solutions to exist [1]. In fact, in systems with purely repulsive particle-particle interactions such as electrons or perpendicularly aligned excitons, this is the mechanism that can drive the solidification [2], and in these repulsive systems quantum Monte Carlo simulations have confirmed the existence of normal solids [15, 13, 14]. When there is a bound state solution, the missing self-interaction terms will open up an energy gap between the particle and anti-particle excitations [10]. The anti-particle excitations are then localized in the self-consistently created wells while the particle excitations with their positive energies are free-running waves. The resulting solid would be incompressible since the potential wells lead to a prohibitive energy cost against doubly occupied sites [10], analogous to a Mott insulator.
In the present study we investigate the possibility of solidification coexisting with quantum coherence in a supersolid. We extend the standard Gross-Pitaevskii formalism to include the effect of the absence of self-interactions specific to this exciton system. In place of the many-body wave function of the normal state system, we construct the order parameter for the incompressible supersolid of the form,
$\Psi(\mathbf{r})=\sqrt{\rho_{B}}+\sum_{i=1}^{N}\phi_{i}(\mathbf{r})\ ,$ (8)
@@
where $i$ is the site index for the $N$ sites. With this construction, each exciton has the same wave function spread across the entire lattice and with peaks at every site. This order parameter results in both off-diagonal long-range order and also diagonal long-range order, the latter coming from the lattice symmetry. In addition, for a supersolid there must exist phase locking between adjacent sites. This arises from mutual overlap of the $\phi_{i}(\mathbf{r})$ [19]. With the $\phi_{i}(\mathbf{r})$ exponentially localized on their sites it is not possible to link the phases on the different lattice sites, since zero-point quantum and thermal fluctuations eliminate coherence between the near-negligible overlaps. We ensure significant overlap by means of the uniform background component $\rho_{B}$ in Eq. (8). In cases when $\rho_{B}$ tends to zero there is negligible overlap, so there can be no supersolid. In cases when the $\phi_{i}(\mathbf{r}-\mathbf{r}_{i})$ reduce to zero, $|\Psi(\mathbf{r})|^{2}\sim\rho_{B}$, and Eq. (8) represents the homogeneous order parameter for a superfluid.
@@
$\!\!\!\!\!\!\frac{\hbar^{2}\nabla^{2}}{2M_{\text{X}}}\Psi(\mathbf{r})+\!\int\!\!d^{2}\mathbf{r}^{\prime}\,t_{\text{XX}}(|\mathbf{r}^{\prime}\!-\!\mathbf{r}|)|\Psi(\mathbf{r}^{\prime})|^{2}\,\Psi(\mathbf{r})$
$-\!\!\int_{\mathbf{r}^{\prime}\in i_{\mathbf{r}}}\!\!\!\!\!\!d^{2}\mathbf{r}^{\prime}t_{\text{XX}}(|\mathbf{r}^{\prime}\!-\!\mathbf{r}|)|\Psi(\mathbf{r}^{\prime})|^{2}(\Psi(\mathbf{r})\!-\!\sqrt{\rho_{B}})=\mu\Psi(\mathbf{r}).$ (9)
@@
The mean-field potential in Eq. (9) differs from the mean-field potential in the standard Gross-Pitaevskii equation for dipolar bosons [21], not only because $V_{\text{XX}}$ has been replaced by $t_{\text{XX}}$ to account for two-body correlations, but also because there is an additional potential term. It is this term, that accounts for the absence of self-interactions and makes supersolid solutions possible.
We solve Eq. (9) by evolving in imaginary time [22], starting from a spatially randomized order parameter. Note throughout the numerical calculation there is no restriction that $\Psi(\mathbf{r})$ is of the form of Eq. (8). Because of the non-local interaction $V_{\text{XX}}(\mathbf{r})$, we evaluate the mean-field potential by means of the convolution theorem [23].
Analogously to the normal solid, to identify the supersolid we fix the lattice positions $\{a_{i}\}$ thereby breaking the translational symmetry ourselves. We choose a square lattice even though in two-dimensions a triangular lattice could have a slightly lower energy. The triangular lattice is computationally very demanding and the issue we are interested in addressing here is the existence of a supersolid. In our static calculation, if the energy of the square lattice supersolid is less than the energy of the superfluid, then so will be the energy of the triangular lattice supersolid. Since the excitons are gate confined within a fixed area of the layers, with one exciton per site the lattice constant $a$ for a square lattice is geometrically related to the exciton density $\rho=1/\pi r_{0}^{2}$ by $a=2r_{0}$.
All simulations were performed on a $256^{2}$ grid covering the sample spatial area. Solutions were taken as converged when the relative change in energy in consecutive time steps decreases below $10^{-8}$. After reaching convergence, we have verified for self-consistency that $\Psi(\mathbf{r})$ indeed has the form of Eq. (8), with $\rho_{B}$ not zero. We focus first on the low-density regime, setting the average interparticle spacing $r_{0}=30$. As length and energy scales, we use the effective Bohr radius $a_{B}=\hbar^{2}\epsilon/e^{2}m_{e}^{*}$ and the effective Rydberg Ry.
Figure 1(a) shows the ground-state order parameter $|\Psi(\mathbf{r})|^{2}$ for small interlayer separation $d=1$. The solution $\Psi(\mathbf{r})$ is homogeneous – corresponding to a superfluid: for such a small $d$, the exciton dipole interactions and the self-consistently calculated self-interaction potential are very weak compared with the kinetic energy, too weak to drive the system to a solid.
@@
Figure 2: (a) Dependence on interlayer separation $d$ of $\rho_{B}/\rho$, the amplitude of the homogeneous background component of $\Psi(\mathbf{r})$, for $r_{0}=30$. (b) Energy eigenvalue of the supersolid state ($S$) and corresponding eigenvalue of the superfluid state (SF), calculated without the self-interaction correction.
@@
Figures 1(c) and 1(d) show that when the interlayer spacing $d$ is further increased, the peak in $\Psi(\mathbf{r})$ on each site narrows and the amplitude $\rho_{B}$ of the homogeneous background component decreases. It may seem at first sight surprising that the lattice constant does not depend on $d$ and the effective interaction strength, but this is because the particle density is fixed by the gate charge and the repulsive interaction only allows one particle per unit cell (in a similar manner the lattice constant of a 2D Wigner crystal is independent of the strength of the Coulomb interaction).
@@
Figure 2(b) compares the supersolid energy eigenvalue $\mu_{SS}$ from Eq. (9), and hence with the broken translational symmetry of the fixed lattice, compared with the superfluid energy eigenvalue $\mu_{SF}$ determined without the subtracted self-interaction term in Eq. (9), and thus without the broken translational symmetry. For $d<6$ the two energy eigenvalues are identical, indicating a stable superfluid, but for $d\gtrsim 6$ the energy eigenvalue $\mu_{SF}$ is higher than $\mu_{SS}$ indicating that the incompressible supersolid becomes the ground state for $d\gtrsim 6$.
@@
Figure 3: The phase diagram at zero temperature, as a function of interlayer separation $d$ and average interparticle spacing $r_{0}$. Dashed red line: superfluid to supersolid transition. Dash-dotted grey line: Superfluid and supersolid condensate collapse into a normal electron-hole liquid [24]. Dashed green line: exciton liquid to normal exciton solid transition taken from Ref. [13]. Dotted black line: $d=r_{0}$. Insets: order parameter profile $|\Psi(\mathbf{r})|^{2}$ at the points indicated by grey dots.
@@
Realistic parameters for exploring the supersolid phase can be extracted from Fig. 3 with knowledge of the effective Bohr radius $a_{B}$ for a specific system. As an example, in a TMD bilayer with an hBN barrier, $a_{B}\sim 1$ nm, the supersolid phase is experimentally accessible with interlayer separations $d\sim 10-20$ nm and densities $\sim 1-5\times 10^{11}$ cm-2.
Reference [14] shows that the Berezinskii-Kosterlitz-Thouless temperature for the normal to superfluid transition and the melting temperature of the normal solid-liquid transition are both of the order of a few mRy. The supersolid melting temperature is also a Berezinskii-Kosterlitz-Thouless transition. Thus we expect the supersolid melting temperature to be of the same order, corresponding in the TMD bilayer to $\sim 2$ K, again readily experimentally accessible.
@@
Within the supersolid phase, as the density is increased the exciton repulsive interaction (Eq. (2)) evolves from dipolar-like to Coulomb-like, making the supersolid soften and eventually melt. Just before melting, the system enters the BEC-BCS crossover regime where screening rapidly becomes strong and the system simultaneously loses coherence and crystal structure. The quantum condensed phases lie almost entirely within the BEC regime where screening is negligible, allowing us to describe the system in terms of a Bose condensate with a macroscopic order parameter. This makes our Gross-Pitaevskii formalism valid across the area of the phase diagram encompassing both the supersolid and superfluid phases.
To observe the exciton supersolid, both spatial periodic order and quantum coherence must be experimentally confirmed. For the spatial order, photoluminescence has been used to detect signatures of crystallization in gate-controlled exciton bilayers [31] and optically pumped exciton bilayers [32]. Scanning Tunneling Microscopy (STM) has been used to directly image a 2D crystal lattice of electrons in a van der Waals heterostructure with sub nanometer resolution through a monolayer graphene gate [33], and this technique should also be applicable to imaging the charge distribution of the top layer in a gated exciton bilayer. Existence of quantum coherence can be demonstrated using photoluminescence interference to detect off-diagonal long-range order [34, 35] and to detect vortices [34, 36].
@@
We thank Filippo Pascucci for very useful discussions. This work was supported by Research Foundation – Flanders (FWO-Vl), with contract number 12A3T24N; by the University Research Fund (BOF), with project numbers 53228, G0A9F25N, G0AIY25N, GOH1122N, G061820N, and G060820N; by the Brazilian Council for Research (CNPq), through the PRONEX/FUNCAP, Universal, and PQ programs; and by the Australian Research Council projects CE170100039, DP210101608 and IL230100072.